Liquid octane CH3CH26CH3 will react with gaseous oxygen O2 to produce gaseous carbon dioxide CO2 and gaseous water H2O. Suppose 90. g of octane is mixed with 145. g of oxygen. Calculate the maximum mass of water that could be produced by the chemical reaction. Be sure your answer has the correct number of significant digits.


1
Expert's answer
2021-03-24T01:08:57-0400

CH3CH26CH3 + 11O2 = 3CO2 + 16H2O

M (CH3CH26CH3) = 68.2862 g/mol

M (O2) = 32 g/mol

M (H2O) = 18 g/mol

n = m / M

n (CH3CH26CH3) = 90 / 68.2862 = 1.32 mol

n (O2) = 145 / 32 = 4.53 mol

According to the equation, 1 mole of octane requires 11 moles of O2. Respectively, 1.32 moles of octane require 1.32 x 11 = 14.52 moles of O2.

Therefore, in this case, O2 is the limiting reactant. With all the available O2 being used, the amount of water to be produced is:

n (H2O) =n (O2) / 11 x 16 = 4.53 / 11 x 16 = 6.59 mol

m (H2O) = n x M = 6.59 x 18 = 118.6 g